OpenStudy (anonymous):

Help!! Suppose you wanted to estimate the average IQ of a community and had randomly selected a sample of 200 residents, with a sample mean IQ of 105. You are informed that the population standard deviation for IQ scores is 15 and have decided to set the alpha level at 0.05. Estimate the average IQ of the population in the community.

OpenStudy (kropot72):

The confidence level equals 100*(1 - alpha)%, or in other words, an alpha of 0.05 indicates a 95 percent confidence level. The population mean lies in the interval \[\bar{x}-1.96\frac{\sigma}{\sqrt{n}}<\mu<\bar{x}+1.96\frac{\sigma}{\sqrt{n}}\] where x-bar is the sample mean, n is the sample size and sigma is the standard deviation. Now you need to plug in the values to find the confidence interval for the population mean.

OpenStudy (kropot72):

@yeame73 Are you there?

OpenStudy (anonymous):

I'm sorry - I had to step away.. Are you still there?

OpenStudy (anonymous):

I think I understand. I was confused because it didn't give a confidence interval percentage but using the alpha makes sense.. Thank you!

OpenStudy (kropot72):

You're welcome :)